OpenStudy (anonymous):

Which is the side length of a cube with a volume of 3375 in3? A. B. C. D.

OpenStudy (anonymous):

will give a fan or meadle

OpenStudy (anonymous):

@paki @jagr2713 @xo_kansasprincess_xo any one

OpenStudy (mathstudent55):

Let's say the cube side is s. \(V = s^3 = 3375\) If \(s^3 = 3375\), then \(\large s = \sqrt[3]{3375} \) Take the cubic root of 3375.

OpenStudy (anonymous):

s = 15 if you don't have access to a calculator.

OpenStudy (anonymous):

thanks i will give you fan and meatle!!

OpenStudy (anonymous):

@robtobey

OpenStudy (yolomcswagginsggg):

Woooooo LoLz

OpenStudy (anonymous):

hi